OpenStudy (anonymous):

Write the recursive formula for the geometric sequence. a1 = -2 a2 = 8 a3 = -32

OpenStudy (anonymous):

A. an = -4 + an-1 B. an = -2 + an-1 C. an = -2 • an-1 D. an = -4 • an-1

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

someone help please

OpenStudy (bahrom7893):

@TuringTest

OpenStudy (anonymous):

do you know?

OpenStudy (bahrom7893):

ohhhh are there subscripts in your options?

OpenStudy (anonymous):

yeah

OpenStudy (bahrom7893):

can u write them properly then? a_n-1 is different from a_(n-1)

OpenStudy (cwrw238):

common ratio r is 8 / -2 = -4

OpenStudy (cwrw238):

you multiply each term by -4 to get the next term

OpenStudy (anonymous):

i dont know how to do this at all

OpenStudy (cwrw238):

you don't add!!

OpenStudy (cwrw238):

to get second term you multiply first by by -4 then you multiply the second term by -4 to get third and so on this is whats meant by recursive